Consider a vector V(24, 71). What will be the length of the orthogonal projection of

this vector along the direction of a unit vector u(0.6, 0.8)?


1
Expert's answer
2021-08-05T06:54:27-0400

The length of the projection is

"\vec{V} \cdot \vec{u}= 24*0.6 +71*0.8 =71.2"